A fishing boat traveled 3 hours against a 6 km/hr current. The return trip took only 2 hours. Find the speed of the boat in still water.
Answer by stanbon(57962)

A fishing boat traveled 3 hours against a 6 km/hr current. The return trip took only 2 hours. Find the speed of the boat in still water.
------------
Against current DATA:
Time = 3 hrs ; rate = b-6 : distance = tr = 3(b-6) = 3b-18 km
-----------
With current DATA:
Time = 2 hrs ; rate = b+6 ; distance = tr = 2(b+6) = 2b+12 km
-----------
EQUATION:
distance with = distance against
2b+12 = 3b-18
b = 30 kph (speed of the boat in quiet water)
==============
Cheers,
Stan H.